OpenStudy (anonymous):

Pls. Help me to solve this :))) f(4) = sqrt x + 5 + 2 pls. answer this and show the solution :)))

OpenStudy (anonymous):

f(4)=sqrt(4+5)+2=3+2=5

OpenStudy (anonymous):

can you do it vertically ??

OpenStudy (anonymous):

\[f(4)=\sqrt{4+5} +2\]\[=\sqrt{9}+2\]\[=3+2\]=5

OpenStudy (anonymous):

THANKS :)))))))))))))))))))))))))))))))))))